a patienr IV is running at 30 drips per minut. The drip chamber is callibrated to deliver 15 drips per mililter. How many mL/hour will the patient receive?
Answer #1

Solution:

The patient is receiving 30drips per minute.

The calibrated drip to deliver is 15drips per mL.

So, If 15 drips deliver in 1 mL

Therefore 30 drips will be deliver = 30 ÷ 15 = 2mL.

So, 2mL is being deliver in 1 minute.

1hour = 60 minutes.

In 1 minute 2mL is being delivered

So, in 60 minutes = 60 × 2 = 120mL/ hour.

The patient will receive 120 mL /hour.